UNDERSTANDING THE METHANOL 5ν_{OH} − 4ν_{OH}/1ν_{CH} IVR FRAGMENTATION ASYMMETRY

\underline{RYAN PEARMAN}, and MARTIN GRUEBELE, Department of Chemistry and Beckman Institute for Advanced Science and Technology, University of Illinois, Urbana, Illinois 61801.

The 1:1 resonance between the 5ν_{OH} overtone and the 4ν_{OH}/1ν_{CH} combination band studied by Rizzo and coworkers exhibits wildly different IVR fragmentation patterns for the two resonant states. Efforts to explain this through hidden symmetries of the wavefunction have met with limited success. We apply an improved version of our vibrational matrix element factorization scheme to a fully anharmonic methanol potential energy surface based on Monte-Carlo sampled ab-initio calculations to predict the IVR lineshapes for rotationless transitions. We find that the different linewidths of the resonant pair can be explained based on fluctuations of the local density of coupled states.